OpenStudy (anonymous):

minimum value of a quadratic is -d/4a or D/4a

hartnn (hartnn):

what are d,D ?

OpenStudy (anonymous):

Discriminant, obviosly

hartnn (hartnn):

both ?

OpenStudy (anonymous):

Now ,yes

hartnn (hartnn):

you mean D/4a or -D/4a ?

hartnn (hartnn):

ok

OpenStudy (anonymous):

\[\huge \left| d \right|=\left| D \right|\]

hartnn (hartnn):

y = ax^2+bx+c y' = 2ax +b y'=0 =2ax+b x= -b/2a plug this in y

hartnn (hartnn):

your a is positive or negative ?

OpenStudy (anonymous):

does it matter

OpenStudy (anonymous):

i mean i am asking y co-oridinate of vertex , = minimum value

hartnn (hartnn):

for y = ax^2 +bx+c min value is c-b^2/4a = D/4a

hartnn (hartnn):

for y = -ax^2 +bx+c min value is -c+b^2/4a =- D/4a

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

D/4a

OpenStudy (anonymous):

my a is positive

hartnn (hartnn):

sorry its reverse

OpenStudy (anonymous):

-D/4a

hartnn (hartnn):

for y = ax^2 +bx+c min value is c-b^2/4a = -D/4a

hartnn (hartnn):

infact, irrespective of the value of a

hartnn (hartnn):

always -D/4a

OpenStudy (anonymous):

that's what i was asking actually for hehe

hartnn (hartnn):

:P